Melanie places $6,000 into an account that earns 4% interest compounded daily. Approximately six months later (180 days), she deposits an additional $3,000 in the same account. What is the value of the account after 1 year? (Assume all years have 365 days. Round your answer to the nearest cent.)

Answer by Theo(13342)

present value is 6000
4% interest compounded daily is equal to .04/365.
180 days later the future value is 6000 * (1 + .04/365) ^ 180 = 6119.524613, at which time she deposits an additional 3000 to get a balance of 9119.524613.
6 months later = 365 - 180 = 185 days later, she will have 9119.5246134 * (1 + .04/365) ^ 185 = 9306.290215 in the account.